40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 2 00:54:08 2017, Koji, Update, ASS, Xarm ASS restoration work 7x
    Reply  Fri Jun 2 01:22:50 2017, gautam, Update, ASS, ASS restoration work ASS_Y_recovery.pngASS_Repairs.zipOLs.png
       Reply  Wed Jun 7 14:11:49 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdfY_OL_calib_check.png
          Reply  Wed Jun 7 17:45:11 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation CoilTFs.pdf
             Reply  Thu Jun 8 02:11:28 2017, gautam, Update, ASS, Y-arm coil driver electronics investigation AnalogCheck.pdfAnalogCheck_zoom.pdfArmCtrl.pdf
                Reply  Mon Jun 12 17:42:39 2017, gautam, Update, ASS, ETMY Oplev Pentek board pulled out 
                   Reply  Wed Jun 14 18:15:06 2017, gautam, Update, ASS, ETMY Oplev restored 
                      Reply  Thu Jun 15 01:56:50 2017, gautam, Update, ASS, ETMY Oplev restored ITMY_OLspec.pdf
                         Reply  Tue Jun 27 16:11:28 2017, gautam, Update, Electronics, Coil whitening ETMYT_1182669422.mp4ITMY_whitening_studies.pdf
Message ID: 13082     Entry time: Tue Jun 27 16:11:28 2017     In reply to: 13064
Author: gautam 
Type: Update 
Category: Electronics 
Subject: Coil whitening 

I got back to trying to engage the coil driver whitening today, the idea being to try and lock the DRMI in a lower noise configuration - from the last time we had the DRMI locked, it was determined that A2L coupling from the OL loops and coil driver noise were dominant from ~10-200Hz. All of this work was done on the Y-arm, while the X-arm CDS situation is being resolved.

To re-cap, every time I tried to do this in the last month or so, the optic would get kicked around. I suspected that the main cause was the insufficient low-pass filtering on the Oplev loops, which was causing the DAC rms to rail when the whitening was turned on. 

I had tried some loop-tweaking by hand of the OL loops without much success last week - today I had a little more success. The existing OL loops are comprised of the following:

  • Differentiator at low frequencies (zero at DC, 2 poles at 300Hz)
  • Resonant gain peaked around 0.6 Hz with a Q of ______ (to be filled in)
  • BR notches 
  • A 2nd order elliptic low pass with 2dB passband ripple and 20dB stopband attenutation

THe elliptic low pass was too shallow. For a first pass at loop shaping today, I checked if the resonant gain filter had any effect on the transmitted power RMS profile - turns out it had negligible effect. So I disabled this filter, replaced the elliptic low pass with a 5th order ELP with 2dB passband ripple and 80dB stopband attenuation. I also adjusted the overall loop gain to have an upper UGF for the OL loops around 2Hz. Looking at the spectrum of one coil output in this configuration (ITMY UL), I determined that the DAC rms was no longer in danger of railing.

However, I was still unable to smoothly engage the de-whitening. The optic again kept getting kicked around each time I tried. So I tried engaging the de-whitening on the ITM with just the local damping loop on, but with the arm locked. This transition was successful, but not smooth. Looking at the transmon spot on the camera, every time I engage the whitening, the spot gets a sizeable kick (I will post a video shortly).  In my ~10 trials this afternoon, the arm is able to stay locked when turning the whitening on, but always loses lock when turning the whitening off. 

The issue here is certainly not the DAC rms railing. I had a brief discussion with Gabriele just now about this, and he suggested checking for some electronic voltage offset between the two paths (de-whitening engaged and bypassed). I also wonder if this has something to do with some latency between the actual analog switching of paths (done by a slow machine) and the fast computation by the real time model? To be investigated.

GV 170628 11pm: I guess this isn't a viable explanation as the de-whitening switching is handled by the one of the BIO cards which is also handled by the fast FEs, so there isn't any question of latency.

With the Oplev loops disengaged, the initial kick given to the optic when engaging the whitening settles down in about a second. Once the ITM was stable again, I was able to turn on both Oplev loops without any problems. I did not investigate the new Oplev loop shape in detail, but compared to the original loop shape, there wasn't a significant difference in the TRY spectrum in this configuration (plot to follow). This remains to be done in a systematic manner. 

Plots to support all of this to follow later in the evening.

Attachment #1: Video of ETMY transmission CCD while engaging whitening. I confirmed that this "glitch" happens while engaging the whitening on the UL channel. This is reminiscent of the Satellite Box glitches seen recently. In that case, the problem was resolved by replacing the high-current buffer in the offending channel. Perhaps something similar is the problem here?

Attachment #2: Summary of the ITMY UL coil output spectra under various conditions.

 

Attachment 1: ETMYT_1182669422.mp4  2.642 MB  Uploaded Wed Jun 28 01:24:19 2017
Attachment 2: ITMY_whitening_studies.pdf  65 kB  Uploaded Wed Jun 28 01:24:39 2017  | Show | Show all
ELOG V3.1.3-